# Bayesian Inference for Consistent Predictions in Overparameterized Nonlinear Regression

### v1

@@ Abstract.
The remarkable generalization performance of overparameterized models has challenged the conventional wisdom of statistical learning theory. While recent theoretical studies have shed light on this behavior in linear models or nonlinear classifiers, a comprehensive understanding of overparameterization in nonlinear regression models remains lacking. This paper explores the predictive properties of overparameterized nonlinear regression within the Bayesian framework, extending the methodology of adaptive prior based on the intrinsic spectral structure of the data. We establish posterior contraction for single-neuron models with Lipschitz continuous activation functions and for generalized linear models, demonstrating that our approach achieves consistent predictions in the overparameterized regime. Moreover, our Bayesian framework allows for uncertainty estimation of the predictions. The proposed method is validated through numerical simulations and a real data application, showcasing its ability to achieve accurate predictions and reliable uncertainty estimates. Our work advances the theoretical understanding of the blessing of overparameterization and offers a principled Bayesian approach for prediction in large nonlinear models.
@@ 1. Introduction
In recent years, the field of machine learning has witnessed a paradigm shift in understanding the generalization performance. Traditional statistical theory emphasized the importance of controlling model complexity and avoiding overfitting by balancing the bias-variance tradeoff [55]. However, the emergence of large-scale deep neural networks, which often have far more parameters than the number of training data, has challenged this conventional wisdom. These models have demonstrated the ability to achieve high accuracy on test data while completely interpolating the training data. This discrepancy between these empirical observations and traditional learning theory has necessitated a new explanation [61].
@@
A groundbreaking study by [5] provided a theoretical foundation in the overparameterized linear regression models, showing the non-asymptotic analysis of the generalization error of a minimum norm interpolator. This discovery has reshaped our understanding of generalization performance in large statistical models, suggesting that overparameterization does not necessarily lead to poor prediction, even in simple models. Instead, the key to good performance lies in the model’s expressiveness and its ability to capture the essential structure of the data. Indeed, subsequent theoretical studies have demonstrated that overparameterization leads to accurate predictions in various statistical models, shedding light on the behavior of modern machine learning models.
@@
While the theoretical understanding of the goodness of overparameterization has advanced significantly in linear estimators, such as linear regression [5, 52] and kernel methods [31, 33], the analysis of nonlinear estimators remains limited. Although some studies have investigated this issue in nonlinear neural networks for classification problems [13, 26], a comprehensive understanding of why and under what conditions overparameterized nonlinear regression can achieve good generalization performance is still lacking. Hence, addressing this question is crucial for reinforcing the theoretical foundations of modern machine learning.
@@
In this paper, we explore the predictive properties of overparameterized nonlinear regression models within the Bayesian framework. Bayesian inference combines the prior distribution of parameters with a likelihood function to obtain the posterior distribution. [59] proposed a methodology that incorporates the structure of the data into the prior distribution while capturing the linear model structure as the likelihood, resulting in accurate predictions even in non-sparse high-dimensional models. The key idea is to adaptively adjust the prior variance based on the intrinsic structure (spectra) of the data, which is found to be critical in prediction error convergence in the sense of posterior contraction. We extend this to nonlinear regression models, specifically single neuron models [60, 47] and generalized linear models (GLMs) [37], and theoretically analyze the prediction error of the resulting posterior distribution. The main contributions of this study are summarized as follows:
@@
  • For generalized linear models based on one-parameter exponential families, including logistic and Poisson regressions, we establish conditions under which the posterior distribution converges to the true parameter in terms of the predictive risk and provide an upper bound on the convergence rate (Theorem 1). This finding highlights the applicability of our approach to the regression of discrete response variables, such as those encountered in classification problems.
@@
  • We show that the proposed method generalizes well in single-neuron models with a Lipschitz continuous nonlinear activation function and Gaussian noise (Theorem 2). This result suggests that, even in more general models involving nonlinearity, overparameterization allows us to predict benignly under appropriate prior distributions.
@@
  • Our Bayesian approach not only provides a theoretical guarantee of the generalization performance of the overparametrized models but also estimates the uncertainty in the predictions. By sampling from the posterior distribution, we can construct a confidence level for each prediction, which is essential in decision-making processes.
@@ 1.1. Related work
Bayesian GLM in high-dimension has been studied intensively. The first paper to study high-dimensional Bayesian GLM [18] achieved posterior concentration for the parameter without assuming any sparsity structure. However, it imposed the restricted growth condition on the parameter dimension, and this underparameterized setting differs from the present work. Subsequently, several studies have explored the GLM setting where the number of parameters may exceed the sample size [24, 34, 23]. These works have demonstrated the asymptotic properties of the posterior distribution, but they assume sparsity, implying that the parameters are inherently low-dimensional. The properties of Bayesian variable selection in GLMs have also been investigated in similar domains [8, 9, 42, 30, 22]. However, the assumption of parameter sparsity can be stringent, and the current study is conducted in a setting that does not rely on this assumption, thus expanding the scope of research in this area. The methods and theory of Bayesian inference in high-dimensional models are summarized in [2].
Numerous theoretical results suggest that overparameterization leads to good generalization performance. [5] demonstrated that the excess risk converges to zero in overparameterized linear models. This theory was widely extended in other linear estimators, such as regularized regression [52, 29, 58, 36, 11], kernel regression [31, 32], linear neural networks [10, 50], and regression for dependent data [40, 53]. [13, 6, 26] have ventured beyond these linear frameworks to investigate the benign nature of overparameterization, but their focus is limited to classification problems. [47] tackled nonlinear regression and demonstrated the absence of benign overfitting, but their assumptions differ from the present work mainly in terms of distributional assumptions on the covariate.
@@ 1.2. Organization
The remainder of this paper is structured as follows. In Section 2, we introduce the problem setting of one-parameter generalized linear models under the overparameterization assumption and provide a detailed description of our proposed methodology. Section 3 presents the theoretical results on posterior contraction for GLMs. Section 4 extends the results to a two-parameter case. We then conduct simulation studies in Section 5 to validate our theoretical findings and assess the uncertainty estimation. Section 6 showcases a real data application, comparing our approach with baseline methods in terms of classification performance. In Section 7, we conclude this paper by discussing the contribution and limitations of this work. The appendix provides the proof and discussion of the main theorems and the details of the computational algorithms used in our analysis.
@@ 1.3. Notation
For $q\in\mathbb{N}$ and any vector $\bm{v}$, $\|\bm{v}\|_{q}$ denotes the $\ell_{q}$-norm of $\bm{v}$. For any vector $\bm{v}$ and symmetric matrix of the same dimension $A$, we define $\|\bm{v}\|_{A}=\sqrt{\bm{v}^{\top}A\bm{v}}$. For sequences $\{a_{n}\}_{n\in\mathbb{N}}$ and $\{b_{n}\}_{n\in\mathbb{N}}$, $a_{n}\gtrsim b_{n}$ denotes the existence of a constant $c>0$ such that $a_{n}\geq cb_{n}$ for all $n\geq\overline{n}$ with some finite $\overline{n}\in\mathbb{N}$, and $a_{n}\lesssim b_{n}$ denotes the opposite. $a_{n}=o(b_{n})$ indicates that $a_{n}/b_{n}\to 0$ as $n\to\infty$ whereas $a_{n}=\omega(b_{n})$ does $a_{n}/b_{n}\to\infty$ as $n\to\infty$. $a_{n}\gtrapprox b_{n}$ denotes that $a_{n}\gtrsim b_{n}(\log n)^{c}$ holds for every $c>0$, and $a_{n}\lessapprox b_{n}$ denotes the opposite. For natural numbers $a,b\in\mathbb{N}$ with $a<b$, we define $[a:b]:=\{a,a+1,a+2,...,b\}$. For any square matrix $A$, let $A^{-1}$ be the inverse matrix of $A$, $A^{\dagger}$ be the Moore–Penrose pseudoinverse matrix, and $\mathrm{tr}(A)$ be the sum of the diagonals of $A$. For a pseudo-metric space $(S,d)$ and a positive value $\delta$, $\mathcal{N}(\delta,S,d)$ denotes the covering number, that is, the minimum number of balls that cover $S$ with a radius $\delta$ in terms of $d$. For two probability measures $\Pi_{1}$ and $\Pi_{2}$ on the same measurable space $(\mathcal{X},\mathcal{B})$, the total variation distance is given by $\|\Pi_{1}-\Pi_{2}\|_{\mathrm{TV}}:=\sup_{B\in\mathcal{B}}|\Pi_{1}(B)-\Pi_{2}(B)|$, and the Kullback–Leibler (KL) divergence and the KL variation (also referred to as the centered second-order KL divergence) between them are respectively defined as $K(\Pi_{1},\Pi_{2})=\int_{\mathcal{X}}\log(\mathrm{d}\Pi_{1}/\mathrm{d}\Pi_{2})\mathrm{d}\Pi_{1}$ and $V(\Pi_{1},\Pi_{2})=\int_{\mathcal{X}}\{\log(\mathrm{d}\Pi_{1}/\mathrm{d}\Pi_{2})-K(\Pi_{1},\Pi_{2})\}^{2}\mathrm{d}\Pi_{1}$ if $\Pi_{1}$ is absolutely continuous with respect to $\Pi_{2}$; otherwise, both are $\infty$.∞ . The Hellinger distance between two probability distributions $\Pi_{1}$ and $\Pi_{2}$ is defined as $H(\Pi_{1},\Pi_{2})=\sqrt{1-\int_{\mathcal{X}}\sqrt{\mathrm{d}\Pi_{1}\mathrm{d}\Pi_{2}}}$.
@@ 2.1. Setting
Consider a pair $(\bm{X},Y)$, where $\bm{X}$ is an $\mathbb{R}^{p}$-valued centered random covariate and $Y$ is an $\mathbb{R}$-valued univariate response. The GLM is defined as follows:
@@
where $\bm{\beta}^{*}\in\mathbb{R}^{p}$ is the true unknown $p$-dimensional regression coefficient, and $g:\mathbb{R}\to\mathbb{R}$ is a known Lipschitz continuous function. In this setting, the response variable $Y$ follows a general distribution belonging to the one-parameter exponential family, with a density function given by:
@@
where $\mu_{x,\bm{\beta}}\in\mathbb{R}$ is the parameter, $a:\mathbb{R}\to\mathbb{R}$ and $b:\mathbb{R}\to\mathbb{R}$ are known continuously differentiable functions, with $a$ having a nonzero derivative. The function $c:\mathbb{R}\to\mathbb{R}$ is chosen such that $L(y;\mu_{x,\bm{\beta}})$ is a proper probability density.
This formulation encompasses several well-known models: Normal linear models: $Y\mid\bm{X}$ follows a $g(\bm{x}^{\top}\bm{\beta}^{*})$-mean Gaussian distribution with fixed variance, and $g$ is the identity function; Logistic regression model: $Y\mid\bm{X}$ follows a Bernoulli distribution with parameter $g(\bm{x}^{\top}\bm{\beta}^{*})$, and $g$ is the logistic function; Poisson regression model: $Y\mid\bm{X}$ follows a Poisson distribution with parameter $g(\bm{x}^{\top}\bm{\beta}^{*})$, and $g$ is the exponential function. Note that we do not restrict the link function $g$ to canonical, providing flexibility in modeling the relationship between the linear predictor and the response variable.
@@
The primary goal of this study is to examine the generalization performance of a predictor trained on the observations $\mathcal{D}$ in the overparameterized domain. We adopt as the predictor the mean $g(\bm{x}^{\top}\bm{\beta})$, with the estimated parameter $\widehat{\bm{\beta}}$ plugged in. Then, the notion of excess risk is introduced to quantify the generalization performance, measuring the difference between the oracle predictor $g(\bm{X}^{\top}\bm{\beta}^{*})$ and the predictor $g(\bm{X}^{\top}\widehat{\bm{\beta}})$:
@@ 2.2. Method
The predictor construction follows a systematic approach: first, the data is divided into two sets; then, a prior distribution of $\bm{\beta}$ is developed based on the empirical spectral information of $\Sigma$ from one set; next, the likelihood is computed in the other set; finally, the posterior distribution of $\bm{\beta}$ is computed based on the prior and likelihood.
To begin, we divide the dataset $\mathcal{D}$ into two equal parts, $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, each containing $n/2$ observations. While the split ratio is arbitrary, it is crucial that it does not depend on the sample size $n$. This data splitting allows us to use $\mathcal{D}_{1}$ for constructing the prior distribution and $\mathcal{D}_{2}$ for evaluating the likelihood function, ensuring that the data used in these two steps are independent.
@@
where $\bar{\bm{x}}\in\mathbb{R}^{p}$ is the sample mean of the covariates from $\mathcal{D}_{1}$. The factor $2/n$ ensures that $\widehat{\Sigma}$ is an unbiased estimator of the population covariance matrix. We then perform a spectral decomposition of $\widehat{\Sigma}$:
@@
This approximation captures the most significant $k$ eigenvalues and eigenvectors of $\widehat{\Sigma}$, effectively reducing the dimensionality of the covariate space while retaining the most informative directions of variation. The choice of $k$ is a tuning parameter; however, to manage computational cost, we fix it in an interval $[L_{\kappa}:U_{\kappa}]$ such that it satisfies Assumption 2.
@@
Although this prior distribution resembles a $p$-variate Gaussian distribution with mean $\bm{0}$ and covariance $\widehat{\Sigma}_{1:k}$, there are some important distinctions. First, the distribution is truncated, with the indicator function restricting the domain to a centered sphere of radius $R$. This truncation is necessary because the low-rank approximation $\widehat{\Sigma}_{1:k}$ is singular, and defining the density function without support restrictions would cause the integrals over $\mathbb{R}^{p}$ to diverge. Second, unlike [42, 23, 54], we impose a non-sparse prior distribution, even though some parameters may be redundant. This choice is motivated by the desire to retain all potentially relevant covariates in the data and to avoid the challenges associated with sparse priors, such as the need to specify sparsity-inducing hyperparameters.
@@
The Bayesian framework allows for the quantification of events of interest, such as the probability of predictions being close to the true values, through the probability measures defined in this manner. Moreover, as will be demonstrated later, the distributional information provided by the posterior is valuable in practice.
Note that we discard the information about the response variable $y_{i}$ in $\mathcal{D}_{1}$ since this subset was only used to compute the sample covariance matrix of the covariates. If dropping this information seems inefficient, one can consider switching the roles of $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ to construct a new estimator (posterior distribution), and averaging the two estimators in a manner similar to the sample splitting scheme [7]. The theoretical results presented later are preserved for this averaging.

### v2

@@ Abstract.
The remarkable generalization performance of large-scale models has been challenging the conventional wisdom of the statistical learning theory. Although recent theoretical studies have shed light on this behavior in linear models and nonlinear classifiers, a comprehensive understanding of overparameterization in nonlinear regression models is still lacking. This study explores the predictive properties of overparameterized nonlinear regression within the Bayesian framework, extending the methodology of the adaptive prior considering the intrinsic spectral structure of the data. Posterior contraction is established for generalized linear and single-neuron models with Lipschitz continuous activation functions, demonstrating the consistency in the predictions of the proposed approach. Moreover, the Bayesian framework enables uncertainty estimation of the predictions. The proposed method was validated via numerical simulations and a real data application, showing its ability to achieve accurate predictions and reliable uncertainty estimates. This work provides a theoretical understanding of the advantages of overparameterization and a principled Bayesian approach to large nonlinear models.
@@ 1. Introduction
In recent years, the field of machine learning has witnessed a paradigm shift in understanding generalization performance. Traditional statistical theory emphasizes the importance of controlling model complexity and avoids overfitting by balancing the bias-variance trade-off [56]. However, with the emergence of large-scale deep neural networks, which often have considerably more parameters than the number of training data samples, conventional wisdom has been challenged. These models have demonstrated the ability to achieve high accuracy on test data by completely interpolating the training data. This discrepancy between empirical observations and traditional learning theory has necessitated a new explanation [62].
@@
A groundbreaking study by [6] provides a theoretical foundation for overparameterized linear regression models through a non-asymptotic analysis of the generalization error of a minimum norm interpolator. This discovery has reshaped our understanding of generalization performance in large statistical models, suggesting that overparameterization does not necessarily lead to poor prediction, even in simple models. Instead, the key to good performance lies in the model’s expressiveness and its ability to capture the essential structure of the data. Recent theoretical studies have demonstrated that overparameterization leads to accurate predictions in various statistical models, shedding light on the behavior of modern machine learning models.
@@
While the theoretical understanding of the goodness of overparameterization has advanced significantly in linear estimation, such as in linear regression [6, 53] and kernel methods [32, 34], the analysis of nonlinear estimators remains limited. Although some studies have investigated this issue in nonlinear neural networks for classification problems [14, 27], a comprehensive understanding is lacking as to why and under what conditions overparameterized nonlinear regression can achieve good generalization performance. Hence, addressing this question is crucial for reinforcing the theoretical foundations of modern machine learning.
@@
In this paper, the predictive properties of overparameterized nonlinear regression models within the Bayesian framework are explored. Bayesian inference combines the prior distribution of parameters with a likelihood function to obtain the posterior distribution. [60] proposed a methodology that incorporates the structure of the data into the prior distribution, capturing the linear model structure as the likelihood, which results in accurate predictions even in non-sparse high-dimensional models. The key idea is to adaptively adjust the prior variance based on the intrinsic structure (spectra) of the data, which is found to be critical in prediction error convergence in the sense of posterior contraction. We extend this to nonlinear regression models, specifically single-neuron models [61, 48] and generalized linear models (GLMs) [38] and theoretically analyze the prediction error of the resulting posterior distribution. The main contributions of this study are summarized as follows:
@@
  • For generalized linear models based on one-parameter exponential families, including logistic and Poisson regressions, the conditions under which the posterior distribution converges to the true parameter are established in terms of predictive risk, providing an upper bound on the convergence rate (Theorem 1). This finding highlights the applicability of the proposed approach to discrete response-variable regression, such as those encountered in classification problems.
@@
  • The proposed method is demonstrated to generalize well to single-neuron models with a Lipschitz continuous nonlinear activation function and Gaussian noise (Theorem 2). This result suggests that even in more general models involving nonlinearity, overparameterization allows benign prediction under appropriate prior distributions.
@@
  • The proposed Bayesian approach not only provides a theoretical guarantee of the generalization performance of the overparametrized models but also estimates the uncertainty in the predictions. By sampling from the posterior distribution, a confidence level can be constructed for each prediction, which is essential in decision-making processes.
@@ 1.1. Related work
Bayesian GLM has been studied intensively for high dimensions. The first paper to study high-dimensional Bayesian GLM [19] realized posterior concentration for the parameter without assuming any sparsity structure by imposing a restricted growth condition on the parameter dimension. This underparameterized setting differs from the present work. Subsequently, several studies explored the GLM setting where the number of parameters may exceed the sample size [25, 35, 24]. These works have demonstrated the asymptotic properties of the posterior distribution; however, they assume sparsity, implying that the parameters are inherently low-dimensional. The properties of Bayesian variable selection in GLMs have also been investigated in similar domains [9, 10, 43, 31, 23]. However, the assumption of parameter sparsity can be stringent, and the current study is conducted in a setting that does not rely on this assumption, thus expanding the scope of research in this area. The methods and theory of Bayesian inference in high-dimensional models are summarized in [2].
Numerous theoretical results suggest that overparameterization leads to good generalization performance. [6] demonstrated that the excess risk converges to zero in overparameterized linear models. This theory has been widely extended to other linear estimators, such as regularized regression [53, 30, 59, 37, 12], kernel regression [32, 33], linear neural networks [11, 51], and regression for dependent data [41, 54]. [14, 7, 27] ventured beyond these linear frameworks to investigate the benign nature of overparameterization, but their focus was limited to classification problems. [48] tackled nonlinear regression and demonstrated the absence of benign overfitting, but their assumptions differ from the present work mainly in terms of distribution of the covariate.
@@ 1.2. Organization
The remainder of this paper is structured as follows. In Section 2, the problem setting for one-parameter generalized linear models under the overparameterization assumption is introduced, providing a detailed description of the proposed methodology. Section 3 presents the theoretical results on posterior contraction for GLMs. Section 4 extends the results to the two-parameter case. Section 5 reports on the simulation studies conducted to validate the theoretical findings and assess the uncertainty estimation. Section 6 showcases a real data application, comparing the proposed approach with baseline methods in terms of classification performance. Section 7 concludes the paper, discussing the contribution and limitations of this work. The proof and discussion of the main theorems and details of the computational algorithms used in the analysis are provided in the Appendix.
@@ 1.3. Notation
For $q\in\mathbb{N}$ and any vector $\bm{v}$, $\|\bm{v}\|_{q}$ denotes the $\ell_{q}$-norm of $\bm{v}$. For any vector $\bm{v}$ and symmetric matrix of the same dimension $A$, we define $\|\bm{v}\|_{A}=\sqrt{\bm{v}^{\top}A\bm{v}}$. For sequences $\{a_{n}\}_{n\in\mathbb{N}}$ and $\{b_{n}\}_{n\in\mathbb{N}}$, $a_{n}\gtrsim b_{n}$ denotes the existence of a constant $c>0$ such that $a_{n}\geq cb_{n}$ for all $n\geq\overline{n}$ with some finite $\overline{n}\in\mathbb{N}$, and $a_{n}\lesssim b_{n}$ denotes the opposite; $a_{n}=o(b_{n})$ indicates that $a_{n}/b_{n}\to 0$ as $n\to\infty$, whereas $a_{n}=\omega(b_{n})$ indicates that $a_{n}/b_{n}\to\infty$ as $n\to\infty$; $a_{n}\gtrapprox b_{n}$ indicates that $a_{n}\gtrsim b_{n}(\log n)^{c}$ holds for every $c>0$, and $a_{n}\lessapprox b_{n}$ denotes the opposite. For natural numbers $a,b\in\mathbb{N}$ with $a<b$, we define $[a:b]:=\{a,a+1,a+2,...,b\}$. For any square matrix $A$, let $A^{-1}$ be the inverse matrix of $A$, $A^{\dagger}$ be the Moore–Penrose pseudoinverse matrix, and $\mathrm{tr}(A)$ be the sum of the diagonals of $A$. For a pseudo-metric space $(S,d)$ and a positive value $\delta$, $\mathcal{N}(\delta,S,d)$ denotes the covering number, that is, the minimum number of balls that cover $S$ with a radius $\delta$ in terms of $d$. For two probability measures $\Pi_{1}$ and $\Pi_{2}$ on the same measurable space $(\mathcal{X},\mathcal{B})$, the total variation distance is given by $\|\Pi_{1}-\Pi_{2}\|_{\mathrm{TV}}:=\sup_{B\in\mathcal{B}}|\Pi_{1}(B)-\Pi_{2}(B)|$, and the Kullback–Leibler (KL) divergence and the KL variation (also referred to as the centered second-order KL divergence) between them are respectively defined as $K(\Pi_{1},\Pi_{2})=\int_{\mathcal{X}}\log(\mathrm{d}\Pi_{1}/\mathrm{d}\Pi_{2})\mathrm{d}\Pi_{1}$ and $V(\Pi_{1},\Pi_{2})=\int_{\mathcal{X}}\{\log(\mathrm{d}\Pi_{1}/\mathrm{d}\Pi_{2})-K(\Pi_{1},\Pi_{2})\}^{2}\mathrm{d}\Pi_{1}$ if $\Pi_{1}$ is absolutely continuous with respect to $\Pi_{2}$; otherwise, both are $\infty$.∞ . The Hellinger distance between two probability distributions $\Pi_{1}$ and $\Pi_{2}$ is defined as $H(\Pi_{1},\Pi_{2})=\sqrt{1-\int_{\mathcal{X}}\sqrt{\mathrm{d}\Pi_{1}\mathrm{d}\Pi_{2}}}$.
@@ 2.1. Setting
Consider the pair $(\bm{X},Y)$, where $\bm{X}$ is an $\mathbb{R}^{p}$-valued centered random covariate, and $Y$ is an $\mathbb{R}$-valued univariate response. The GLM is defined as follows:
@@
where $\bm{\beta}^{*}\in\mathbb{R}^{p}$ is the true unknown $p$-dimensional regression coefficient, and $g:\mathbb{R}\to\mathbb{R}$ is a known Lipschitz continuous function. In this setting, the response variable $Y$ follows a general distribution belonging to the one-parameter exponential family with a density function given by
@@
where $\mu_{x,\bm{\beta}}\in\mathbb{R}$ is the parameter; $a:\mathbb{R}\to\mathbb{R}$ and $b:\mathbb{R}\to\mathbb{R}$ are known continuously differentiable functions, with $a$ having a nonzero derivative. The function $c:\mathbb{R}\to\mathbb{R}$ is chosen such that $L(y;\mu_{x,\bm{\beta}})$ is a proper probability density.
This formulation encompasses several well-known models: Normal linear models: $Y\mid\bm{X}$ follows a $g(\bm{x}^{\top}\bm{\beta}^{*})$-mean Gaussian distribution with fixed variance, and $g$ is the identity function; Logistic regression model: $Y\mid\bm{X}$ follows a Bernoulli distribution with parameter $g(\bm{x}^{\top}\bm{\beta}^{*})$, and $g$ is the logistic function; Poisson regression model: $Y\mid\bm{X}$ follows a Poisson distribution with parameter $g(\bm{x}^{\top}\bm{\beta}^{*})$, and $g$ is the exponential function. Note that the link function $g$ is not restricted to being canonical, thereby providing flexibility in modeling the relationship between the linear predictor and response variable.
@@
The primary objective of this study is to examine the generalization performance of a predictor trained on the observations $\mathcal{D}$ in the overparameterized domain. We adopt as the predictor the mean $g(\bm{x}^{\top}\bm{\beta})$, with the estimated parameter $\widehat{\bm{\beta}}$ plugged in. Then, the notion of excess risk is introduced to quantify the generalization performance, measuring the difference between the oracle predictor $g(\bm{X}^{\top}\bm{\beta}^{*})$ and the predictor $g(\bm{X}^{\top}\widehat{\bm{\beta}})$:
@@ 2.2. Method
Predictor construction follows a systematic approach: first, the data are divided into two sets; then, a prior distribution of $\bm{\beta}$ is developed based on the empirical spectral information of $\Sigma$ from one set; next, the likelihood is computed in the other set; finally, the posterior distribution of $\bm{\beta}$ is computed based on the prior and likelihood.
Initially, we divide the dataset $\mathcal{D}$ into two equal parts, $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, each containing $n/2$ observations. Although the split ratio is arbitrary, it should not depend on the sample size $n$. The data splitting allows the use of $\mathcal{D}_{1}$ for constructing the prior distribution and $\mathcal{D}_{2}$ for evaluating the likelihood function, ensuring that the data used in these two steps are independent.
@@
where $\bar{\bm{x}}\in\mathbb{R}^{p}$ is the sample mean of the covariates from $\mathcal{D}_{1}$. The factor $2/n$ ensures that $\widehat{\Sigma}$ is an unbiased estimator of the population covariance matrix, whereby a spectral decomposition of $\widehat{\Sigma}$ is performed:
@@
This approximation captures the most significant $k$ eigenvalues and eigenvectors of $\widehat{\Sigma}$, effectively reducing the dimensionality of the covariate space while retaining the most informative directions of variation. In choosing $k$ as a tuning parameter to manage the computational cost, it has to be fixed to an interval of $[L_{\kappa}:U_{\kappa}]$ so that it satisfies Assumption 2.
@@
Although this prior distribution resembles a $p$-variate Gaussian distribution with mean $\bm{0}$ and covariance $\widehat{\Sigma}_{1:k}$, there are some important distinctions. First, the distribution is truncated, with the indicator function restricting the domain to a centered sphere of radius $R$. This truncation is necessary because the low-rank approximation $\widehat{\Sigma}_{1:k}$ is singular, and defining the density function without support restrictions would cause the integrals over $\mathbb{R}^{p}$ to diverge. Second, unlike [43, 24, 55], we impose a non-sparse prior distribution, even though some parameters may be redundant. This choice is motivated by the desire to retain all potentially relevant covariates in the data and avoid the challenges associated with sparse priors, such as the need to specify sparsity-inducing hyperparameters.
@@
The Bayesian framework allows the quantification of events of interest, such as the probability of predictions being close to the true values through probability measures defined in this manner. Moreover, as demonstrated later, the distributional information provided by the posterior is valuable in practice.
Note that the information about the response variable $y_{i}$ in $\mathcal{D}_{1}$ is discarded since this subset is only used to compute the sample covariance matrix of the covariates. If excluding this information seems inefficient, the roles of $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ can be switched to construct a new estimator (posterior distribution), and the two estimators can be averaged in a manner similar to the sample splitting scheme [8]. The theoretical results presented next preserve this averaging.